OpenStudy (anonymous):

Write the slope-intercept equation for the line with the given slope that contains the given point. m = 8/3; (3, −4)

OpenStudy (anonymous):

I started with y=8/3(3) + ?

OpenStudy (slaaibak):

\[y - y_1 = m(x - x_1)\] Use that equation

OpenStudy (slaaibak):

where (3.-4) = (x1, y1)

OpenStudy (anonymous):

\[y - (-4) = \frac{8}{3}(x - 3)\] and solve for y

OpenStudy (anonymous):

ohhh ok thank you...then what do I do after that please?

OpenStudy (anonymous):

Just solve for y and you're done.

OpenStudy (slaaibak):

you have to have it in this form: y = mx + C

OpenStudy (anonymous):

ok thanx so much

OpenStudy (anonymous):

@ Tyler & slaaibak